From the 1/14th Daily Journal Summaries for the month of March, 1968.
The Daily Journal Summaries
1 March 1968
"Golden Dragons" continued show of force operations in the Bong Son Plain conducting local patrols by 2 rifle companies and the reconnaissance platoon. Fire base construction continues.
Plans: Continue show of force operations.
2 March 1968
B/1-50th Mech requested a Dust-Off for 1 US WIA from sniper fire vic Bridge 91(921182). Dust-Off complete at 1319H. Enemy fire came from vic 917165, est 2-3 AW.
1/B/1-50th Mech vic 917165 detained 2 VN. Indivs were trying to exit area of sniper fire. Both had ID cards and were evac to B/1-50th Mech CP.
"Golden Dragons" continued operations in PATRICK AO on a reduced scale. Show of force operations were conducted by Delta Company. Maximum effort was placed on personnel and services for the rifle companies at LZ GERONIMO. B/1-50th (Mech) swept the area where sniper fire had been received along highway #1 with negative findings.
Plans: The security company at LZ GERONIMO will be changed. Charlie Company will be relieved by Alpha Company. Charlie Company will cdt operations to the North searching for NVA waystation locations.
3 March 1968
B/1-50th Mech detained 2 indivs vic of bridges and evac to LZ LO BOY. Detainees were questioned and released.
BS874105; 1 NVA Bn, neg reports of wpns.
BS905076; 1 Co, Unknown type. (Both C-3).
Subject: Possible Enemy Country-Wide Offensive
In view of the fact that 3 March 68 is the anniversary of the founding of the LAO DONG PARTY and other indications of impending enemy activity in the I and II Corps Tactical Zones; a widespread coordinated country-wide attack will occur within the next 24-36H.
1-14th Infantry continued refit and resupply of all elements prior to cdt of offensive operations. Negative sightings or contacts were reported during this period.
Plans: Delta Company relieves elements of 1-35 at LZ LARAMIE providing scty for C/6-16 Arty(155). Company C moves to the northwest conducting search and destroy operations. Alpha Compnay prov scty LZ GERONIMO and cdts local sweeps. B Company remains in LZ ENGLISH (OPCON 1-35). B/1-50(Mech) reverts to parent unit.
4 March 1968
"Golden Dragons" conducted operations within AO PATRICK with one rifle company and relieved elements of 1-35 at LZ LARAMIE. Negative sightings of contact was reported in the AO. B/1-50th(Mech) returned to LZ UPLIFT and control of 1-50th Mech.
Plans: "Golden Dragons" will continue operations in Bong Son Plain and begin planning in AN LOA Valley.
5 March 1968
1-14th Infantry cdts operations within the Bong Son Plain. North of the Bon Son River, secures LZ LARAMIE with one rifle company and LZ GERONIMO with one rifle company. Local sweeps were conducted around both fire bases by Delta and Alpha Companies respectively. Negative sightings of contacts were reported.
Plans: Cdt operations in AN LOA Valley with CAs on LZ TAPE w/Delta Co(-) and Hvy Mortar Platoon, CAs in the northern protion of AN LOA with Charlie Company.
6 March 1968
1/D/1/14 vic 738134 observed movement to their front and engaged resulting in 1 VC KIA approx 17-20 yrs old.
"Golden Dragons" conducted offensive operations in the AN LOA Valley. Delta Company(-) combat assaulted to LZ TAPE provided security for Hvy Mortar Platoon. Charlie Company conducted a combat assault to the northern portion of the AN LOA.
Delta Company made contact w/one VC vicinity LZ TAPE. Alpha Company observed muzzle flashes north of LZ GERONIMO.
Plans: Continue operations in AN LOA with Charlie Company moving from night location to the South.
7 March 1968
"Golden Dragons" continued operations in the An Loa Valley below the "Fish Hook" with one company conducting search and destroy operations. Delta Co(-) secured Hvy Mortar Platoon and conducted local sweeps around LZ TAPE.
Plans: Continue operations in the An Loa, conduct combined operation with 40th ARBN east of LZ TOM and conduct local sweep around LZ GERONIMO.
8 March 1968
"Golden Dragons" conducted a combined operation with the 40th ARVN in the coastal plain east of LZ TOM. Charlie Company was extracted from the AL LOA Valley and inserted to the Northwest of LZ SANTANA to sweep the infiltration route north and south of LZ SANTANA. Delta Company secured Hvy Mortar Platoon until it was extracted to LZ LARAMIE and 2 tubes into LZ GERONIMO.
Plans: Continue stay behind tactics in the AN LOA VALLEY with Delta Company-sweep infiltration routes vic LZ SANTANA.
9 March 1968
D/1/14 vic 744098 had 1 VC KIA and 1 VC WIA CIA. Indivs tried to evade and were engaged. Dust-Off requested for VC WIA CIA and complete at 1425H. Also CIA 3 packs containing rice, greens, candy, etc. Destroyed.
1930H: Bde S2 reports that the VC CIA WIA was a CPL in the 2nd VC Regt.
The "Golden Dragons" continued operations in the AN LOA Valley and Bong Son Plain. Delta Company made contact with food carriers south of LZ TAPE.
Plans: Continue operations in the AN LOA with Delta Company and exchange Bravo Company with Charlie Company at LZ ENGLISH. Conduct an operation with C/1-10th Cav in order to seek out the sniper along Highway #1.
10 March 1968
A/1/14 and 3/C/1-10th Cav closed together. Loc 920166 3/C/1-10th Cav detained 7 indivs that tried to run from U.S. troops. 1 Indiv was evac to Bde; others released.
Intell Report: MI TOR 1150H (F-6) BS875125: 2 NVA companies in this vic wearing green uniforms and training the local people in use of wpns.
3/A/1/14 w/ARVN Intell Team vic 924165 had contact with 4-5 VC in spider holes resulting in 2 VC KIA dressed in grey and 1 S&W Cal .38 CIA.
3/D/1/14 vic 750047 spotted 1 VC and engaged resulting in 1 VC KIA wearing black PJ's; 18 yrs old and carrying a pack containing 2 sets of clothes.
"Golden Dragons" continued offensive operations in PATRICK AO conducting operations in the AN LOA Valley and also in the Bong Son Plain. Alpha Company(-) conducted a combat assualt to the beach with the ARVN Inetll Squad from the 40th ARVN. 3/C/1-10th Cav blocked and Alpha Company swept the village of Chau Me for the sniper harassing the engineers along Highway #1.
Delta Company(-) continued S&D opns in the AN LOA. Charlie Company continued OPCON 1-35 at LZ ENGLISH.
Plans: Continue security at LZ GERONIMO, LARAMIE and LZ ENGLISH and continue conducting operations in the AN LOA Valley against ride carrying parties.
11 March 1968
Rcn/1/14 requested a Dust-Off for 6 US WIA from a dud M-79 rd vic of LZ GERONIMO. Dust-Off requested and complete at 1115H.
2/Rcn/1/14 loc 778188 spotted enemy carying weapons and engaged w/SA fire resulting in 1 VC KIA and 2 VC WIA CIA and 1 SKS CIA SN642115. Also spotted 2 more enemy w/wpns moving toward them. Neg assessment.
Fox Force made contact with VC with weapons on the trail complex vic Hill 529. Delta observed movement in the AN LOA valley and engaged the target with Tango Deuce. Results unknown.
Plans: Continue operations in the AN LOA Valley and conduct a combined operation in the Bong Son Plain with C/1-10th Cav(-) and Alpha Company(-) and an ARVN Intell Squad.
12 March 1968
A/1/14(-) vic 872122 detained 6 VN Males (5 of military age). Evac to Bde M.I.
"Golden Dragons" continued operations conducting a combat assault with Alpha Company(-) to cordon and Search the village of Tan AN(1). Results were 6 detainees. Operations continued in the AN LOA Valley with Bravo Company and Delta(-). The Delta 10 was lifted from LZ ENGLISH (48 Hr Stand down) to the AN LOA Valley.
Plans: Extract Fox Force from the AO and continue saturation patrols and ambushes in the AN LOA.
13 March 1968
Rcn/1/14(2) vic 727187 had 3 VC KIA w/packs and 2 NVA pistol belts w/US Canteen and cup. Also had some M-60 ammo in a US first aid pack. Military aged Montagnard.
"Golden Dragons" continued operations in AO PATRICK with 3 rifle compnaies. "Golden Dragons" continued security at LZ GERONIMO & LARAMIE. Charlie Company continued OPCON 1-35 at LZ ENGLISH.
Recon made contact with 3 VC and resulted in 3 VC KIA. Delta Company flip flopped the 2nd platoon and 3rd platoon from LZ LARAMIE.
Plans: continue operations in the AN LOA and Bong Son Plain and begin operations aimed at the routes of infiltration vic NUI CA LANG Mts.
14 March 1968
"Golden Dragons" continued operations in PATRICK AO with three(3) rifle companies and a reconnaissance platoon. Charlie Company continued OPCON 1-35 for security at LZ ENGLISH. No significant contact or sightings were reported. Delta Company secured LZ LARAMIE with two platoons. Alpha Company conducted local sweeps vic LZ GERONIMO with two rifle platoons.
Plans: Conduct MEDCAP vic LZ GERONIMO. Insert Fox 1 and conduct combat assault vic MY BINH to search for communications wire reported along SUOI PHU MY RIVER.
15 March 1968
Intell Report: Rating F6: 1 NVA Company(100 men) loc in block 882140-890140-885145-885138. Dressed in black uniforms w/SA. Moved into the area 141800H.
1-14th Infantry continued operations in the AN LOA VALLEY and Alpha Company(-) conducted a combat assault vic MY BINH to search for communications lines. C/1-10th Cav(-) linked up with Alpha Company(-). Alpha Company detained 3 VCS. Fox 3 was inserted via Charlie Charlie vic LZ SANDRA.
Plans: Prepare to execute Phase II and relieve elements C/1-10th Cav as bridge security force. Delta and Alpha Company will flip flop.
16 March 1968
"Golden Dragons" continued operations in the Northern portion of PATRICK AO and relieved elements of C/1-10th Cav on Bridges 87-94. Delta Company assumed security along highway #1. Alpha Company assumed security for LZ LARAMIE. Bravo continued operations in AN LOA VALLEY.
Plans: Continue operations in the AN LOA Valley conducting night ambushes and night patrols vic LZ GERONIMO. Plans to conduct CHECKMATE OPN are being made.
17 March 1968
D/1/14(-) detained 3 indivs vic of GERONIMO and evac to LZ ENGLISH at 1540H. 1 indiv had anti-US propaganda.
The "Golden Dragons" conducted operations in PATRICK AO with negative contact reported. Delta Company continued security at LZ GERONIMO and provided dcty for Bridges 87-94 along highway #1. Bravo Company continued saturation patrols in the AN LOA Valley. Alpha Company continued S&D opns on infiltration routes north of LZ LARAMIE.
Battalion conducted a "SNATCH" operation in the village of PHU NONG which resulted in three VCS, one of which was a VC Cadreman.
Plans: 1-14th Infantry continues operations directed at the VC infrastructure in the Bong Son Plain with an ARVN Intell Squad. Delta Company will conduct a CHECKMATE operation on highway #1.
18 March 1968
2(-), CP/D/1/14 w/Intell Squad had 1 VCS female vic 869176. Evac to Bde.
"Golden Dragons" continued night operations in the AN LOA Valley with Bravo Company. Delta Company worked the local hamlets with the VC Cadreman picked up in the "SNATCH" OPN on 17 March 1968. Fox 1 was inserted and 3 was extracted. Delta Company detained 1 VCS female and found a box of M-79 rds in a hedgerow. An ARVN Intell Squad worked with Delta Company.
Plans: Bravo Company will be extracted from the AN LOA Valley and inserted in the Bong Son Plain and sweep local villages and conduct night operations. Delta Company will conduct local sweep with the Hamlet Chief to point out Vietcong in the area.
19 March 1968
1 US KIA (Engr) 0830H 300 meters north of Bridge 89. Indiv killed by a mine in the road that he was detonating. (881188)
1/D/1/14 vic 867185 detained 1 VC (Female). Indiv was a VC resupply woman and her husband is a NVA Official that was captured by the ARVN.
"Golden Dragons" concluded operations in the AN LOA VALLEY with Bravo Company combat assaulting into the highground overlooking the Bong Son Plain. Delta Company continued local sweeps around LZ GERONIMO and conducted a CHECKMATE OPERATION at Bridge #88 with no unusual results reported.
Plans: Bravo Company will sweep into the Bong Son Plain while Delta Company will establish blocking positions south of LZ GERONIMO.
20 March 1968
4/D/1/14 vic Bridge 88 reported that a VN Civilian hit what was believed to be an anti-personnel mine along the side of the road. One leg badly mangled.
1-14th Infantry continued operations in the Bong Son Plain with Bravo, Delta(-) and Fox 1. The operation concluded with negative results and an obvious lack of movement in the area.
Delta Company reported a VN civilian hit an anti-personnel mine in the vic of Bridge 88. Delta Company vic Bridge88 rec'd approximately 30 rds AW fire from an AK-47. Negative casualties were reported and fire was returned with M-60 and .50 Cal.
Plans: Bravo Company will conduct movement to the West of highway #1 to over watch the highway activity. Delta Company will continue local patrolling vic LZ GERONIMO. Alpha company continues security LZ LARAMIE and conducts S&D operations on infiltration routes North of LZ LARAMIE. Charlie Company will continue security LZ ENGLISH, OPCON 1-35.
21 March 1968
0530H: D/1/14 sub el vic bridge 94 reported 4 81mm rds landed vic 919214. 15 VN WIA and evac to Sa Huynh. Rds were friendly H&I fire from LZ CHARLIE BROWN.
"Golden Dragons" continued operations in the Bong Son Plain with emphasis on the high ground overlooking highway #1. Delta Company continued security along Highway #1 and received 4 rds 81mm vic Bridge 94. 15 VN WIA and 1 VN KIA resulted as two rds landed in a village east of the highway. Bravo Company continued moving to overlook Bridges 93 and 88. Fox 1 eas extracted to LZ GERONIMO. Alpha Company(-) continued S&D opns to the North of LZ LARAMIE, Charlie Company remains OPCON 1-35 at LZ ENGLISH.
Plans: Bravo Company continues to saturate the area west of highway #1 to seek out the sniper along highway #1. Delta Company will conduct a CA in combined operation with 3rd 40th ARVN.
22 March 1968
1/B/1/14 vic 907199 had 1 VC KIA; 1 VC Detainee and 1 SKS CIA SN7059308. Found 1 B40 rocket case vic 908197.
1435H: 1/B/1/14 had 3 Male and 1 Female detainee military age. Found misc ammo in hooches.
Intell Report: Bde NCS reports strong indication of widespread attacks at 230300H.
1 Bn 14th Inf continued operations in the Bon Song Plain with Bravo Company working the high ground to the west of LZ GERONIMO overlooking hwy #1. In vic 908197 1/B/1/14 made contact with 3 armed VC resulting in 1 VC KIA, 1 VC CIA and 1 SKS CIA. Delta Company(-) conducted a CA in support of operations of the 3/40th ARVN with neg results. Mort/1/14 Inf fired on 3 VC in the AN LOA Valley but were unable to make an assessment of damage due to darkness. Alpha Company continued S&D opns in the high ground north of LZ LARAMIE. Charlie Company remains OPCON 1-35 at LZ ENGLISH.
Plans: Continue S&D opns in assigned AO. Bravo Company to conduct cordon and search opns in vicinity LZ WILLIE(7895).
23 March 1968
D/1/14 reported a 3/4 ton truck rec'd approx 5-6 snoper fire between Bridges 84 and 83. 1 US WIA(Slight).
1-14th Infantry continued operations in PATRICK AO with 3 rifle companies. Hvy mortar platoon observed movement in the AN LOA Valley and placed fire on them with unknown results. Fox 2 was inserted to act as a stay behind in the Bravo 10 location. A 3/4 ton vehicle from Delta Company received 5-6 sniper rds vic Bridge 83.
Plans: Cordon and Search operation vicinity LZ WILLIE (7895) was delayed 24 hrs due to aircraft assets. Delta and Alpha Company will conduct flip flops.
24 March 1968
"Golden Dragons" conducted operations in the Bong Son Plain with Bravo Company working the high ground east of LZ GERONIMO. Mort/1/14 engaged several VC w/packs moving in the AN LOA Valley; results unknown. Alpha Company continued S&D opns north of LZ LARAMIE.
Plans: Bravo Company to conduct cordon and search operations vic LZ WILLIE, Delta Company will conduct local S&D vic LZ GERONIMO. Alpha Company continues S&D opns along possible infiltration routes north of LZ LARAMIE.
25 March 1968
B/1/14 vic 790952 w/Intell Squad had 23 indivs to be evac to Bde S2. Mission complete at this time.
1 Bn 14 Infantry continued operations in PATRICK AO. Bravo Company conduct a cordon and search operations in vic LZ WILLIE which resulted in 23 VC suspects captured. Delta Company conducted local S&D opns from LZ GERONIMO with neg results. Rcn/1/14 reported hearing 2 rifle shots south of LZ GERONIMO but a sweep of the area resulted in negative findings. Mort/1/14 reported 1 VC w/pack in AN LOA Valley moving north. Alpha Company continued to sweep the area north of LZ LARAMIE.
Plans: Continue S&D opns in the assigned AO with Bravo Company moving into the high ground north of LZ WILLIE. Elements of the Recon Plt are to be inserted to the northwest of LZ GERONIMO.
26 March 1968
3/D/1/14 vic 871177 picked up a VCS whose son is a hamlet guerilla. Evac to Bde.
"Golden Dragons" continued opns in assigned AO. At approximately 0300 LZ GERONIMO (Bn Fire Base) received 15-20 rounds of small arms ammo. The firing caused no US casualties, enemy losses are unknown. A Company detected movement outside the perimeter (vic 801063) but there was no contact. A and B Companies continued S&D opns north of LZ WILLIE (vic 7896) with neg results. Mort/1/14 observed 5 VC moving north (vic753036), engaged with unknown results.
Plans: Continue S&D opns in assigned AO with A and B Companies working the high ground north of LZ WILLIE. D Company will continue local sweeps off of LZ GERONIMO.
27 March 1968
"Golden Dragons" continued operations in PATRICK AO. In the early morning hours MINI BASE FOX (vic LZ GERONIMO) received one incoming grenade with negative casualties. Illum was called for with negative findings. A and B Companies continued to work the high ground north of LZ WILLIE. B Co engaged 3 VC w/packs and wpns. Area sweep resulted in negative findings. D Company conducted local sweeps north of LZ GERONIMO with negative results.
Plans: Continue S&D opns in southern portion of AO. B Company to move by air to assume responsibility for security of the Bong Son Bridge.
28 March 1968
3/D/1/14 vic 880164 had 1 detainee (male) w/possible important information. Evac to LZ GERONIMO.
B/1/14 vic Bon Song Bridge had 1 Us indiv drown after diving into the river at that location. It was believed that he suffered from cramps or struck his head. Dustoff was caled and indiv was DOA. Time: 1145H.
1/14 continued operations in assigned AO. B Co was air lifted to the Bon Song Bridge at 0915. A Co continued to operate in the high ground north of LZ WILLIE with negative findings. D Company while conducting local sweeps south of LZ GERONIMO picked up 1 male VC suspect. C Company continued to provide security for LZ ENGLISH.
Plans: Continue S&D operations in AO with A Company moving west into the southern end of the An Loa Valley. B Company to secure th Bon Song Bridge. Delta Company conduct local sweeps from LZ GERONIMO. C Company continues security at LZ ENGLISH.
29 March 1968
1/14 continued operations in PATRICK AO. A Company moved west into the An Loa valley and came under sniper fire. A Company called in arty and gunships. Contact was broken at 1730H with negative US casualties and unknown enemy damage assessment. B Company continued to secure the Bong Son Bridge. C Company continued securtity at LZ ENGLISH. D Company conducted local sweeps north of LZ GERONIMO.
Plans: A Company to continue to work north in the An Loa. D Company to conduct local sweeps from LZ GERONIMO with one platoon. Bridges will be secured by 3/C/1-10 and Delta 30/40 will be lifted to LZ GERONIMO.
31 March 1968
3/C/1-10th Cav requested a Dustoff for an ARVN soldier WIA from a sniper vic 920185. Dustoff requested and complete at 1210H.
3/C/1-10th Cav vic 920185 detained 1 military aged male w/ID Card that had been tampered with. Evac to LZ GERONIMO.